Online concert tickets cost $37 each plus a service charge of $8.50 per ticket the web site also charges a transaction fee of $14.99 for the purchase you paid $242.49 how many tickets did you buy
Answer by richwmiller(17219)

(37+8.50)*t+14.99=242.49
45.50*t=242.49-14.99
45.50*t=227.5
t=227.5/45.50
t=5